# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ Abstract.
Exponential actions defined by vector configurations provide a universal framework for several constructions of holomorphic dynamics, non-Kähler complex geometry, toric geometry and topology. These include leaf spaces of holomorphic foliations, intersections of real and Hermitian quadrics, the quotient construction of toric varieties, LVM and LVMB manifolds, complex-analytic structures on moment-angle manifolds and their partial quotients. In all of these cases, the geometry and topology of the appropriate quotient object can be described by combinatorial data including a pair of Gale dual vector configurations.
@@
The work was supported by the Russian Science Foundation under grant no. 23-11-00143, http://rscf.ru/en/project/23-11-00143/, and performed at Steklov Mathematical Institute of Russian Academy of Sciences.
@@ 1. Introduction and notation
There is a remarkable connection between the linear properties of the configuration $\Gamma$ and the topology of the foliation $\mathbb{R}^{m}$ by the orbits of the action (1.1). This connection is most effectively described by the Gale dual configuration of vectors $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in the space $W^{*}\cong\mathbb{R}^{m-k}$.
@@
The key topological fact is that the action of $V$ on $U(\mathcal{K})$ is free and proper if and only if the data $\{\mathcal{K};{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ defines a simplicial fan $\Sigma$ in the space $W^{*}$. Using Gale duality, the fan criterion can be formulated directly in terms of the original configuration $\Gamma$ defining the action. Moreover, the orbit space $U(\mathcal{K})/V$ is compact if and only if the fan $\Sigma$ is complete. In this case $U(\mathcal{K})/V$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}$, and for the holomorphic action on $\mathbb{C}^{m}$ with the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}=(D^{2},S^{1})^{\mathcal{K}}$. Finally, if the configuration $\Gamma$ (or, equivalently, $\mathrm{A}$) lies in a lattice, then the fan $\Sigma$ is rational, and the action (1.1) extends to an algebraic torus action on $U(\mathcal{K})\subset\mathbb{C}^{m}$ whose quotient space is the toric variety $X_{\Sigma}$.
Also of interest is the class of exponential actions for which the corresponding fan $\Sigma$ is the normal fan of a convex polytope. There is a criterion for a fan to be normal in terms of the Gale dual cones. (In the rational case, this gives a criterion for the toric variety $X_{\Sigma}$ to be projective.) When $\Sigma$ is a normal fan, the orbit space $U(\mathcal{K})/V$ is identified with the nondegenerate intersection of real quadrics (and in the case of a holomorphic action, with the intersection of Hermitian quadrics).
In Section 2 we identify free orbits of the action (1.1) or, equivalently, nondegenerate leaves of the foliation of $\mathbb{R}^{m}$ by $V$-orbits. A point ${\textbf{{x}}}\in\mathbb{R}^{m}$ has a nontrivial stabiliser if and only if the subcofiguration of $\Gamma$ corresponding to the set of nonzero coordinates of x does not span $V^{*}$. The set of nonfree orbits is therefore an arrangement of coordinate planes in $\mathbb{R}^{m}$ corresponding to the universal simplicial complex $\mathcal{K}(\Gamma)=\{I\subset[m]\colon\Gamma_{\widehat{I}}\text{ spans }V^{*}\}$ associated with $\Gamma$. Any subcomplex $\mathcal{K}\subset\mathcal{K}(\Gamma)$ defines an open subset in $\mathbb{R}^{m}$, namely,
@@
The topology of the space of free $V$-orbits can be quite bad unless we consider subsets $U(\mathcal{K})\subset\mathbb{R}^{m}$ such that the the restriction of the $V$-action (1.1) to $U(\mathcal{K})$ is proper. In this case, the space of orbits $U(\mathcal{K})/V$ is Hausdorff and has a smooth atlas of charts transverse to $V$-orbits. It is this quotient manifold $U(\mathcal{K})/V$ that features in the constructions of algebraic geometry and topology mentioned above. The necessary and sufficient condition for the $V$-action on $U(\mathcal{K})$ to be proper is the fan condition, which is most readily formulated in terms of the Gale dual configuration.
@@
We set up the notation for linear Gale duality in Section 3. From now on, we work with a pair of Gale dual configurations $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ and $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$. The ‘fan condition’ (Section 4) is that the data $\{\mathcal{K},\mathrm{A}\}$ defines a (simplicial) fan $\Sigma$. The latter means the the cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}=\mathbb{R}_{\geqslant}\langle{\textbf{{a}}}_{i}\colon i\in I\rangle$ corresponing to $I\in\mathcal{K}$ do not overlap pairwise, that is, the relative interior of $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ does not intersect the relative interior of $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ for different $I,J$ in $\mathcal{K}$. In terms of the Gale dual configuration $\Gamma$, this is equivalent to the following: the Gale dual cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$ and $\mathop{\mathrm{cone}}\Gamma_{\widehat{J}}$ do overlap for $I,J$ in $\mathcal{K}$ (Theorem 4.3, Theorem 4.4). This criterion appeared in several works on combinatorial convexity, often implicitly, and it is in the heart of the duality between fans and bunches developed in [1, §2.2].
@@
In Section 5 we prove that the exponential action of $V$ on $U(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a fan data (Theorem 5.1). This property is of purely topological nature. On the other hand, it is the algebraic and holomorphic versions of this result that have become well known. The ‘fan condition’ features in the quotient construction of toric varieties, it guarantees that the quotient variety is separated, i. e. Hausdorff in the usual topology, see [15, Theorem 5.1.11]. The Gale dual to the fan condition appeared as the ‘imbrication condition’ in the work [9] as part of the definition of a class of non-Kähler complex manifolds arising as the leaf spaces of holomorphic foliations, now known as LVMB manifolds. The fact that the ‘fan condition’ and the ‘imbrication condition’ are equivalent via Gale duality was used implicitly in several subsequent works relating complex geometry of LVMB manifolds to algebraic geometry of toric varieties [28, 35, 8, 7]. In the topological setting, the relationship between the fan condition and the separatedness of the quotient featured in the works [22] and [21].
@@
In Section 7 the quotient $U(\mathcal{K})/V$ corresponding to a complete simplicial fan $\Sigma$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$ or the polyhedral product $(D^{1},S^{0})^{\mathcal{K}}$. Thus, the extensive information available on the topology of moment-angle manifolds, including the known descriptions of their cohomology rings and homotopy type, can be brought to bear on the study the topology of foliations by orbits of exponential actions. Topology of moment-angle manifolds and polyhedral products is currently a very active area of research, there is an extensive literature on this subject, see [13, 3].
Normal fans of polytopes and their associated presentations of the quotients $U(\mathcal{K})/V$ by intersections of quadrics are the subject of Section 8. The criterion for a fan $\Sigma=\{\mathop{\mathrm{cone}}A_{I}\colon I\in\mathcal{C}\}$ to be the normal fan of a polytope in terms of the Gale dual configuration $\Gamma$ is stated in Theorem 8.3: instead of intersecting pairwise, the relative interiors of all Gale dual cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$, $I\in\mathcal{C}$, must have a common intersection. This criterion plays an important role in algebraic geometry of toric varieties [29] and holomorphic dynamics and complex geometry of LVM manifolds [26, 27]. There are two ways to relate (the normal fans of) polytopes to intersections of quadrics, which are Gale dual to each other.
@@
and produces an intersection of quadrics defined by the Gale dual data of $(\mathrm{A},{\textbf{{b}}})$:
@@
The intersection of quadrics is nondegenerate (and therefore smooth) if and only if the polytope is generic (in particular, simple) [13, Theorem 6.1.3]. The nondegenerate intersection of quadrics is identified with the quotient $U(\mathcal{K})/V$ by the proper exponential action defined by the normal fan of $P$ (Theorem 8.4). The proof mimics the reconstruction of a Hamiltonian toric manifold from its moment polytope via symplectic reduction [16, 2, 20]; the intersection of quadrics (the moment-angle manifold), appears here as the nondegenerate level set for the moment map. The relationship with the original symplectic construction is reviewed in §14.3.
@@
There is an important particular case when the vectors of the $\mathrm{A}$-configuration satisfy a linear relation with all positive coefficients. In this case, $\mathop{\mathrm{cone}}A$ is the whole of $W^{*}$, while $\mathop{\mathrm{cone}}\Gamma$ is a strongly convex (pointed) cone. The Gale dual vector configuration $\Gamma$ can be replaced by an affine point configuration $\Gamma^{\prime}$ in an affine hyperplane of $V^{*}$ (the affine Gale transform of $\mathrm{A}$). The ‘fan condition’ guaranteeing the properness of the exponential action can be restated in terms of the convex hulls of points $\mathop{\mathrm{conv}}\Gamma^{\prime}_{\widehat{I}}$ instead of cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$. This is precisely the setup used in the definition of LVM and LVMB manifolds, see Section 9.
@@
When the configuration $\Gamma$ is rational, that is, its $\mathbb{Z}$-span $L^{*}=\mathbb{Z}\langle\Gamma\rangle$ is a discrete subgroup (a lattice) in $V^{*}$, both the real and holomorphic exponential actions extend to an action of the algebraic torus $\mathbb{C}^{\times}_{L}=L\otimes_{\mathbb{Z}}\mathbb{C}^{\times}\cong V\times T_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$, where $T_{L}=L\times_{\mathbb{Z}}S^{1}$ is the compact torus defined by $L$. If the algebraic action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper, then the data $\{\mathcal{K},\mathrm{A}\}$ defines a fan $\Sigma$ which is rational with respect to the Gale dual lattice $N=\mathbb{Z}\langle\mathrm{A}\rangle$ in $W^{*}$. The quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ is the toric variety $V_{\Sigma}$ defined by the rational fan $\Sigma$. We obtain a holomorphic principal bundle $\mathcal{Z}_{\mathcal{K}}\to V_{\Sigma}$ with fibre a holomorphic compact torus $\mathbb{C}^{\times}_{L}/\widetilde{V}\cong T_{L}$, known as the generalised Calabi–Eckmann fibration. Perturbing the vector configuration $\Gamma$ (or $\mathrm{A}$) destroys its rationality, the closed holomorphic tori in the fibres of the bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ ‘open up’, and the fibre bundle turns into a holomorphic foliation $\mathcal{F}$ of $\mathcal{Z}_{\mathcal{K}}$ with noncompact leaves. Holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ model irrational deformations of toric varieties [5, 7, 33, 24, 23, 6].
Thinking of $V$ as the space of linear relations between the vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$, we observe that $\Gamma$ is a rational configuration if and only if $V$ is generated by relations with integer coefficients. More generally, a subspace $Q\subset V$ is rational (with respect to $\Gamma$) if it is generated by relations with integer coefficients (Contruction 11.2). A rational subspace $Q\subset V$ contains a full-rank lattice
@@
The exponential action of $V$ on $\mathbb{R}^{m}$ extends to an action of the group $\mathbb{C}^{\times}_{L}\times_{Q}V\cong T_{L}\times V$ on $\mathbb{C}^{m}$, called the torus-expential action (Section 12). Once again, the torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{Q}V)\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a fan data (Theorem 12.2). The quotient $U_{\mathbb{C}}(\mathcal{K})/(T_{L}\times V)$ by a proper torus-exponential action is identified with the quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$ and is called the partial quotient of a moment-angle manifold. There is also a holomorphic version of torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{\widetilde{Q}}{\widetilde{V}})\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$, defined by a rational complex subspace $\widetilde{Q}\subset\widetilde{V}$. If $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, this defines a complex-analytic structure on the partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$ as the quotient $U_{\mathbb{C}}(\mathcal{K})/(\mathbb{C}^{\times}_{L}\times_{\widetilde{Q}}{\widetilde{V}})$, with a holomorphic torus principal bundle $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/T_{L}$. Complex moment-angle manifolds $\mathcal{Z}_{\mathcal{K}}$ are universal examples of complex manifolds with maximal torus action. According to a classification result of [21], any such manifold is biholomorphic to a holomorphic partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$. Two cases of holomorphic partial quotients $\mathcal{Z}_{\mathcal{K}}/T_{L}$, corresponding to a one-dimensional rational subspace $Q\subset V$ and the whole of $V$ being rational, respectively, are of particular importance.
@@
LMB manifolds are a class of non-Kähler complex-analytic manifolds introduced by Lopez de Medrano and Verjovsky [26] and Meersseman [27] as the leaf spaces of exponential holomorphic foliations on the projectivisations of open subsets $U_{\mathbb{C}}(\mathcal{K})\subset\mathbb{C}^{m}$. This construction was further generalised by Bosio [9] and became known as LVMB manifolds. In the original construction of [9] the conditions ensuring that the leaf space is Hausdorff (or the action is proper) were formulated in terms of overlapping convex regions (see Section 13). These Bosio’s conditions are equivalent to the fan condition on the Gale dual data $\{\mathcal{K},\mathrm{A}\}$ (Proposition 13). This allows one to identify LVMB manifolds with holomorphic partial quotients of $\mathcal{Z}_{\mathcal{K}}$ by the diagonal circle, corresponding to a one-dimensional rational subspace $Q\subset V$. In these terms, LVM manifolds can be described as LVMB manifolds corresponding to the normal fans of polytopes, and they are diffeomorphic to nondegenerate intersections of special Hermitian quadrics.
@@
Toric manifolds or smooth complete toric varieties $X_{\Sigma}$ are partial quotients of moment-angle manifolds $\mathcal{Z}_{\mathcal{K}}$ by a maximal freely acting subtorus. The corresponding holomorphic torus principal bundles $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ are described in Section 14. Theorem 8.3 gives a criterion for the rational fan $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ of $X_{\Sigma}$ to be the normal fan of a polytope in terms of the Gale dual cones, which is equivalent to projectivity of $X_{\Sigma}$. This criterion can be extended to a description of the nef and ample cones of $X_{\Sigma}$ ([1, Proposition 2.4.2.6] and Proposition 14.1):
$\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})=\bigcap_{I\in\mathcal{C}}\mathop{\mathrm{cone}}\Gamma_{\widehat{I}},\qquad\mathop{\mathrm{missing}}{Ample}(X_{\Sigma})=\bigcap_{I\in\mathcal{C}}\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}.$
@@
  • $P$ a convex polyhedron in $W^{*}$. A polytope is a bouned polyhedron.
@@
  • $N=M^{*}$ the full rank lattice spanned by a rational vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$.

### v2

@@ Abstract.
Exponential actions defined by vector configurations provide a universal framework for several constructions of holomorphic dynamics, non-Kähler complex geometry, toric geometry and topology. These include leaf spaces of holomorphic foliations, intersections of real and Hermitian quadrics, the quotient construction of simplicial toric varieties, LVM and LVMB manifolds, complex-analytic structures on moment-angle manifolds and their partial quotients, reviewed in this survey. In all cases, the geometry and topology of the appropriate quotient object can be described by combinatorial data including a pair of Gale dual vector configurations.
@@
The work was funded within the framework of the HSE University Basic Research Program
@@ 1. Introduction and notation
There is a remarkable connection between the linear properties of the configuration $\Gamma$ and the topology of the foliation in $\mathbb{R}^{m}$ by the orbits of the action (1.1), which is the subject of this survey. The connection is most effectively described by the Gale dual configuration of vectors $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in the space $W^{*}\cong\mathbb{R}^{m-k}$.
@@
The key topological fact is that the action of $V$ on $U(\mathcal{K})$ is free and proper if and only if the data $\{\mathcal{K};{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ define a simplicial fan $\Sigma$ in the space $W^{*}$. Using Gale duality, the fan criterion can be formulated directly in terms of the original configuration $\Gamma$ defining the action. Moreover, the orbit space $U(\mathcal{K})/V$ is compact if and only if the fan $\Sigma$ is complete. In this case $U(\mathcal{K})/V$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}$, and for the holomorphic action on $\mathbb{C}^{m}$ with the moment-angle manifold $\mathcal{Z}_{\mathcal{K}}=(D^{2},S^{1})^{\mathcal{K}}$. Finally, if the configuration $\Gamma$ generates a lattice (a discrete subgroup) in $V^{*}$, then the fan $\Sigma$ is rational, and the action (1.1) extends to an algebraic torus action on $U(\mathcal{K})\subset\mathbb{C}^{m}$ whose quotient space is the toric variety $X_{\Sigma}$.
Also of interest is the class of exponential actions for which the corresponding fan $\Sigma$ is the normal fan of a convex polytope. There is a criterion for a fan to be normal in terms of the Gale dual cones. (In the rational case, this gives a criterion for the toric variety $X_{\Sigma}$ to be projective.) When $\Sigma$ is a normal fan, the orbit space $U(\mathcal{K})/V$ is identified with the nondegenerate intersection of real quadrics and, in the case of a holomorphic action, with the intersection of Hermitian quadrics.
In Section 2 we describe free orbits of the action (1.1) or, equivalently, nondegenerate leaves of the foliation of $\mathbb{R}^{m}$ by $V$-orbits. The stabiliser of a point ${\textbf{{x}}}\in\mathbb{R}^{m}$ is trivial if and only if the subconfiguration of $\Gamma$ corresponding to the set of nonzero coordinates of x spans $V^{*}$. The set of nonfree orbits is therefore an arrangement of coordinate planes in $\mathbb{R}^{m}$ corresponding to the universal simplicial complex $\mathcal{K}(\Gamma)=\{I\subset[m]\colon\Gamma_{\widehat{I}}\text{ spans }V^{*}\}$ associated with $\Gamma$. Any subcomplex $\mathcal{K}\subset\mathcal{K}(\Gamma)$ defines an open subset in $\mathbb{R}^{m}$, namely,
@@
The topology of the space of free $V$-orbits can be quite bad unless we consider subsets $U(\mathcal{K})\subset\mathbb{R}^{m}$ such that the restriction of the $V$-action (1.1) to $U(\mathcal{K})$ is proper. In this case, the space of orbits $U(\mathcal{K})/V$ is Hausdorff and has a smooth atlas of charts transverse to $V$-orbits. It is this quotient manifold $U(\mathcal{K})/V$ that features in the constructions of algebraic geometry and topology mentioned above. The necessary and sufficient condition for the $V$-action on $U(\mathcal{K})$ to be proper is the fan condition, which is most readily formulated in terms of the Gale dual configuration.
We set up the notation for linear Gale duality in Section 3. From now on, we work with a pair of Gale dual configurations $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ in $V^{*}$ and $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$. The ‘fan condition’ (Section 4) is that the data $\{\mathcal{K},\mathrm{A}\}$ define a (simplicial) fan $\Sigma$. A pair $\{\mathcal{K},\mathrm{A}\}$ satisfying the fan condition is known as a triangulated vector configuration [18], and this concept was brought into the context of holomorphic foliations and LVMB manifolds by Battaglia and Zaffran in [7].
@@
The fan condition means that the cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}=\mathbb{R}_{\geqslant}\langle{\textbf{{a}}}_{i}\colon i\in I\rangle$ corresponding to $I\in\mathcal{K}$ do not overlap pairwise, that is, the relative interior of $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ does not intersect the relative interior of $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ for different $I,J$ in $\mathcal{K}$. In terms of the Gale dual configuration $\Gamma$, this is equivalent to the following: the Gale dual cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$ and $\mathop{\mathrm{cone}}\Gamma_{\widehat{J}}$ do overlap for $I,J$ in $\mathcal{K}$ (Theorem 4.3, Theorem 4.4). This criterion appeared in several works on combinatorial convexity, often implicitly, and it is in the heart of the duality between fans and bunches developed in [1, §2.2].
@@
In Section 5 we prove that the exponential action of $V$ on $U(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a fan data (Theorem 5.1). This property is of purely topological nature. On the other hand, it is the algebraic and holomorphic versions of this result that have become well known. The ‘fan condition’ features in the quotient construction of toric varieties, it guarantees that the quotient variety is separated, i. e. Hausdorff in the usual topology, see [16, Theorem 5.1.11]. The Gale dual to the fan condition appeared as the ‘imbrication condition’ in the work of Bosio [10] as part of the definition of a class of non-Kähler complex manifolds arising as the leaf spaces of holomorphic foliations, now known as LVMB manifolds. The fact that the ‘fan condition’ and the ‘imbrication condition’ are equivalent via Gale duality was used in several subsequent works relating complex geometry of LVMB manifolds to algebraic geometry of toric varieties [35, 43, 9, 7, 8]. In the topological setting, the relationship between the fan condition and the separatedness of the quotient featured in the works [27] and [26].
@@
In Section 7 the quotient $U(\mathcal{K})/V$ corresponding to a complete simplicial fan $\Sigma$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$ or the polyhedral product $(D^{1},S^{0})^{\mathcal{K}}$. Thus, the extensive information available on the topology of moment-angle manifolds, including the known descriptions of their cohomology rings and homotopy type, can be brought to bear on the study of the topology of foliations by orbits of exponential actions. Topology of moment-angle manifolds and polyhedral products is currently a very active area of research, there is an extensive literature on this subject, see [14, 3].
Normal fans of polytopes and their associated presentations of the quotients $U(\mathcal{K})/V$ by intersections of quadrics are the subject of Section 8. The criterion for a fan $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ to be the normal fan of a polytope in terms of the Gale dual configuration $\Gamma$ is stated in Theorem 8.3: instead of intersecting pairwise, the relative interiors of all Gale dual cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$, $I\in\mathcal{C}$, must have a common intersection. This criterion plays an important role in algebraic geometry of toric varieties [36] and holomorphic dynamics and complex geometry of LVM manifolds [33, 34]. There are two ways to relate (the normal fans of) polytopes to intersections of quadrics, which are Gale dual to each other.
@@
and produces an intersection of quadrics defined by the Gale dual data:
@@
The intersection of quadrics is nondegenerate (and therefore smooth) if and only if the polytope is generic (in particular, simple) [14, Theorem 6.1.3]. The nondegenerate intersection of quadrics is identified with the quotient $U(\mathcal{K})/V$ by the proper exponential action defined by the normal fan of $P$ (Theorem 8.4). The proof mimics the reconstruction of a Hamiltonian toric manifold from its moment polytope via symplectic reduction [19, 2, 23]; the intersection of quadrics (the moment-angle manifold), appears here as the nondegenerate level set for the moment map.
@@
There is an important particular case when the vectors of the $\mathrm{A}$-configuration satisfy a linear relation with all positive coefficients. In this case, $\mathop{\mathrm{cone}}\mathrm{A}$ is the whole of $W^{*}$, while $\mathop{\mathrm{cone}}\Gamma$ is a strongly convex (pointed) cone. The Gale dual vector configuration $\Gamma$ can be replaced by an affine point configuration $\Gamma^{\prime}$ in an affine hyperplane of $V^{*}$ (the affine Gale transform of $\mathrm{A}$). The ‘fan condition’ guaranteeing the properness of the exponential action can be restated in terms of the convex hulls of points $\mathop{\mathrm{conv}}\Gamma^{\prime}_{\widehat{I}}$ instead of cones $\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$. This is precisely the setup used in the definition of LVM and LVMB manifolds, see Section 9.
@@
When the configuration $\Gamma$ is rational, that is, its $\mathbb{Z}$-span $L^{*}=\mathbb{Z}\langle\Gamma\rangle$ is a discrete subgroup (a lattice) in $V^{*}$, both the real and holomorphic exponential actions extend to an action of the algebraic torus $\mathbb{C}^{\times}_{L}=L\otimes_{\mathbb{Z}}\mathbb{C}^{\times}\cong V\times T_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$, where $T_{L}=L\times_{\mathbb{Z}}S^{1}$ is the compact torus defined by $L$. If the algebraic action of $\mathbb{C}^{\times}_{L}$ on $U_{\mathbb{C}}(\mathcal{K})$ is proper, then the data $\{\mathcal{K},\mathrm{A}\}$ define a fan $\Sigma$ which is rational with respect to the Gale dual lattice $N=\mathbb{Z}\langle\mathrm{A}\rangle$ in $W^{*}$. The quotient $U_{\mathbb{C}}(\mathcal{K})/\mathbb{C}^{\times}_{L}$ is the toric variety $V_{\Sigma}$ defined by the rational fan $\Sigma$. We obtain a holomorphic principal bundle $\mathcal{Z}_{\mathcal{K}}\to V_{\Sigma}$ with fibre a holomorphic compact torus $\mathbb{C}^{\times}_{L}/\widetilde{V}\cong T_{L}$, known as the generalised Calabi–Eckmann fibration [32, 33, 35]. Perturbing the vector configuration $\Gamma$ destroys its rationality, the closed holomorphic tori in the fibres of the bundle $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ ‘open up’, and the fibre bundle turns into a holomorphic foliation $\mathcal{F}$ of $\mathcal{Z}_{\mathcal{K}}$ with noncompact leaves. Holomorphic foliated manifolds $(\mathcal{Z}_{\mathcal{K}},\mathcal{F})$ model irrational deformations of toric varieties [5, 7, 41, 29, 28, 6].
Thinking of $V$ as the space of linear relations between the vectors ${\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}$, we observe that $\Gamma$ is a rational configuration if and only if $V$ is generated by relations with integer coefficients. More generally, a subspace $Q\subset V$ is rational (with respect to $\Gamma$) if it is generated by relations with integer coefficients (Construction 11.2). A rational subspace $Q\subset V$ contains a full-rank lattice
@@
The exponential action of $V$ on $\mathbb{R}^{m}$ extends to an action of the group $\mathbb{C}^{\times}_{L}\times_{Q}V\cong T_{L}\times V$ on $\mathbb{C}^{m}$, called the torus-exponential action (Section 12). Once again, the torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{Q}V)\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a fan data (Theorem 12.2). The quotient $U_{\mathbb{C}}(\mathcal{K})/(T_{L}\times V)$ by a proper torus-exponential action is identified with the quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$ and is called a partial quotient of a moment-angle manifold. There is also a holomorphic version of torus-exponential action $(\mathbb{C}^{\times}_{L}\times_{\widetilde{Q}}{\widetilde{V}})\times U_{\mathbb{C}}(\mathcal{K})\to U_{\mathbb{C}}(\mathcal{K})$, defined by a rational complex subspace $\widetilde{Q}\subset\widetilde{V}$. If $\{\mathcal{K},\mathrm{A}\}$ is a fan data, this defines a complex-analytic structure on the partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$ as the quotient $U_{\mathbb{C}}(\mathcal{K})/(\mathbb{C}^{\times}_{L}\times_{\widetilde{Q}}{\widetilde{V}})$, with a holomorphic torus principal bundle $\mathcal{Z}_{\mathcal{K}}\to\mathcal{Z}_{\mathcal{K}}/T_{L}$.
Complex moment-angle manifolds $\mathcal{Z}_{\mathcal{K}}$ are universal examples of complex manifolds with maximal torus action. According to a classification result of Ishida [26], any such manifold is biholomorphic to a holomorphic partial quotient $\mathcal{Z}_{\mathcal{K}}/T_{L}$. Two cases of holomorphic partial quotients $\mathcal{Z}_{\mathcal{K}}/T_{L}$, corresponding to a one-dimensional rational subspace $Q\subset V$ and the whole of $V$ being rational, respectively, are of particular importance.
@@
LVM manifolds are a class of non-Kähler complex-analytic manifolds introduced by Lopez de Medrano and Verjovsky [33] and Meersseman [34] as the leaf spaces of exponential holomorphic foliations on the projectivisations of open subsets $U_{\mathbb{C}}(\mathcal{K})\subset\mathbb{C}^{m}$. This construction was further generalised by Bosio [10] and became known as LVMB manifolds. In the original construction of [10] the conditions ensuring that the leaf space is Hausdorff (or the action is proper) were formulated in terms of overlapping convex regions (see Section 13). These conditions by Bosio are equivalent to the fan condition on the Gale dual data $\{\mathcal{K},\mathrm{A}\}$ (Proposition 13.2).
LVMB manifolds are therefore identified with projectivised moment-angle manifolds, that is, holomorphic partial quotients of $\mathcal{Z}_{\mathcal{K}}$ by the diagonal circle (corresponding to a one-dimensional rational subspace $Q\subset V$). In these terms, LVM manifolds can be described as LVMB manifolds corresponding to the normal fans of polytopes, and are diffeomorphic to nondegenerate intersections of special Hermitian quadrics.
@@
Toric manifolds or smooth complete toric varieties $X_{\Sigma}$ are partial quotients of moment-angle manifolds $\mathcal{Z}_{\mathcal{K}}$ by a maximal freely acting subtorus. The corresponding holomorphic torus principal bundles $\mathcal{Z}_{\mathcal{K}}\to X_{\Sigma}$ are described in Section 14. Theorem 8.3 gives a criterion for the rational fan $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ of $X_{\Sigma}$ to be the normal fan of a polytope in terms of the Gale dual cones, which is equivalent to projectivity of $X_{\Sigma}$. This criterion can be extended to a description of the nef and ample cones of $X_{\Sigma}$ ([1, Proposition 2.4.2.6] and Proposition 14.2). We expect a similar criterion to hold in the irrational setting, thereby providing a description of nef and ample divisors on irrational deformations of toric varieties. We finish by describing the construction of Hamiltonian toric manifolds via symplectic reduction and their connection with projective nonsingular toric varieties (Theorem 14.5). Likewise, this connection extends to the irrational setting.
@@
  • $P$ a convex polyhedron in $W^{*}$. A polytope is a bounded polyhedron.
@@
  • $N=M^{*}$ the full rank lattice spanned by or containing a rational vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$.
@@
The author is grateful to the anonymous referees for very helpful comments and suggestions.